OpenStudy (anonymous):

(radical7 +radical2)x(radical7 -3radical2)

OpenStudy (lgbasallote):

\[(\sqrt 7 + \sqrt 2) \times (\sqrt 7 - 3\sqrt 2)\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

use FOIL..just like your usual algebra thingies

OpenStudy (anonymous):

Distribute (or FOIL as some call it). Example: \[\sqrt{3}*\sqrt{3}=\sqrt{9}=3\]

OpenStudy (anonymous):

\[2\sqrt{5}*\sqrt{45}=2\sqrt{225}=2*15=30\]

OpenStudy (anonymous):

i nenver learned how to multiply with radicals o.O

OpenStudy (anonymous):

It's just like I explained above.

OpenStudy (anonymous):

(a + b)(c - d) = (a)(c) - (a)(d) + (b)(c) - (b)(d)

OpenStudy (anonymous):

\[a=\sqrt{7}{\space}{\space}{\space}b=\sqrt{2}{\space}{\space}{\space}c=\sqrt{7}{\space}{\space}{\space}d=3\sqrt{2}\]

OpenStudy (anonymous):

EDIT: d = \[-3\sqrt{2}\]